# Vulnerability Disclosure or Notification? Best Practices for Reaching Stakeholders at Scale

### v1

@@ Abstract.
There has also been research into known vulnerabilities, using datasets or active scans to discover how many machines are still vulnerable. The ethical guidelines suggest that researchers also make an effort to notify the owners of these machines. We posit that this differs from vulnerability disclosure, but rather the practice of vulnerability notification. This practice has some similarities with vulnerability disclosure but should be distinguished from it, providing other challenges and requiring a different approach.
@@ 1. Introduction
Coordinated Vulnerability Disclosure (CVD), formerly known as responsible disclosure, is a best practice in the security community for dealing with discovered vulnerabilities. When a new vulnerability is found, the finder conducts a risk and impact assessment and then initiates conversations with stakeholders to mitigate vulnerable systems or services in time. However, due to the shift in the network landscape, the scale of vulnerable systems and stakeholders involved has changed drastically from the past (van der Horst, 2023; Durumeric et al., 2014). The amount and variety of vulnerabilities have increased (Nakajima et al., 2019), and so have the affected parties (Boucher and Anderson, 2022; Li et al., 2016). This brings a new challenge for security researchers: figuring out a scalable method to identify reliable contact information and ensure the message is delivered to the stakeholders.
@@
Despite the challenges in vulnerability disclosure, notifications to stakeholders with known vulnerabilities that remain in existing systems have gained attention over the years (van der Horst, 2023). We identify the practice as vulnerability notification that informs end-users, such as domain owners, hosting providers, network operators, and incident responders. These stakeholders are different from vendors, who are traditionally the stakeholders in CVD. Whether or not the vulnerability is possibly known to the stakeholders and the public is the difference between vulnerability disclosure and vulnerability notification. A finder discovers a new vulnerability and plans to disclose it to the vendor in vulnerability disclosure, while a finder locates a known vulnerability on existing machines and notifies the end-users in vulnerability notification. This means that vulnerability notification is often conducted after the vulnerability disclosure. In particular, after a newly found vulnerability has been disclosed and mitigated with stakeholders or even made public, it is not guaranteed that all vulnerable systems are treated on a timely basis (Durumeric et al., 2014). Hence, vulnerability notification is there to improve the safety of the internet.
@@
Vulnerability notification is initiated after network scanning (van der Ham and van Rijswijk-Deij, 2017) or vulnerability analysis on existing datasets (Chen et al., 2024) to locate the vulnerable systems with targeted vulnerabilities. A finder should retrieve the stakeholders’ contact information and select a communication channel to reach the affected parties. However, similar to the challenge in vulnerability disclosure, notification to stakeholders also suffers from the high number and complexity of stakeholders (Stivala et al., 2024; Çetin et al., 2017a). The contact retrieval and notification at scale pose an even tougher challenge to finders (Maass et al., 2021; Stock et al., 2016a) since stakeholders have a higher tendency to have multiple parties involved (Çetin et al., 2017a; Chen et al., 2024). Moreover, unlike vulnerability disclosure with CVD as the best practice, best practices for vulnerability notification at scale are still lacking in the community.
@@ 2.1. Vulnerability Disclosure and Notification
Vulnerability disclosure, especially coordinated vulnerability disclosure, has been an accepted practice in security research for years. There have been numerous publications and even standards published on this process (oec, 2021). However, notifying users with vulnerable systems has usually not been seen as a separate practice, but can come with challenges of its own. The difference between the two processes is illustrated in Figure 1.
In vulnerability disclosure, a finder discovers a new vulnerability and aims to share this with the vendor. The role of the finder in the following content includes both the responsibilities of identifying a vulnerability and informing the responsible parties (Forum of Incident Response and Security Teams (2020), FIRST). Once the disclosure process has started, the vulnerability will usually be revealed for the first time. The finder and vendor(s) discuss the finding with or without a coordinator’s support. Subsequently, the vendor works on a mitigation. Vulnerability disclosure needs trusted channels and message details to minimise disclosure leaks to unwanted receivers. The finder may also need to expect extra steps to communicate with the stakeholder on whether or not and how to disclose the new vulnerabilities to the public, based on the disclosure policy and legal agreement on both ends. After some time, the vulnerability is made public, usually with a possible mitigation available.
In vulnerability notification, the vulnerability information is already available to the stakeholders before a notification starts. This means the vulnerability has already been documented and revealed to the stakeholders or the public. The weakness is likely documented with CVE numbers or discussed within specific communities. Potential malicious parties, such as criminals, may also know the existence of the vulnerabilities and can try to abuse them. The receiving users may or may not be aware of the existence of the vulnerability before seeing the notification message. Once they are aware of the issues, they may be able to find the vulnerability information with resources that are not limited to the finder. This can lead to different user response behaviours to the finder. The affected user may also have a different risk assessment based on the severity of the vulnerabilities. The aim of the notification process is to improve general security by removing vulnerable systems.
@@ 2.2. Evaluations of Disclosure and Notification
Vulnerability disclosure has received best practices over the years through the contributions from security researchers, vendors, and support organisations. There have been several publications describing the practices in detail and examining their effectiveness. Householder et al. (Householder and Spring, 2022) proposed a model to assess the coordinator role in CVD, which provides insight into incident handling of vendors and the vulnerability lifecycle. Walshe et al. (Walshe and Simpson, 2022b) examined the effectiveness of how vendors with CVD programs and outsourced vulnerability platforms receive and process vulnerability disclosure. They revealed that disclosure program operators face a large number of vulnerability reports, which becomes a burden in vulnerability prioritisation. Nakajima et al. (Nakajima et al., 2019) looked into vulnerability management of IoT vendors across two countries and pointed out the disclosure pitfalls to avoid.
Vulnerability notification has received growing attention over the last decade, especially in the review of the notification efficiency. The Dutch Institute for Vulnerability Disclosure (DIVD) proposed the notification guideline (van der Horst, 2023) based on their framework and the Communication-Human Information Processing (C-HIP) model (Wogalter et al., 1999) with their notification operation experiences. Their contribution focuses on scalable notification strategies for end-users, such as an incident responder and an abuse specialist. Additionally, the challenges they documented with end-users resemble the capacity and awareness issues of program operators discussed in (Walshe and Simpson, 2022b).
@@ 2.3. Large-Scale Disclosure and Notification
Large-scale disclosure operations have been constantly reported as an increasing challenge by studies in the recent decade (van der Horst, 2023; Householder and Spring, 2022; Walshe and Simpson, 2022b; Nakajima et al., 2019; Moura and Heidemann, 2023; Van Hove et al., 2023; Boucher and Anderson, 2022). Unlike the one-to-one or one-to-multiple disclosure to vendors, the complexity and amount of stakeholders to inform have increased to a level beyond the effort individuals and small teams can make.
Many of the challenges associated with multi-party disclosure processes have been identified by CERT/CC (CERT/CC, 2024) and the FIRST Special Interest Group on Vulnerability Coordination, resulting in a best practice document ”Guidelines and practices for Multi-Party Vulnerability Coordination and Disclosure”(Forum of Incident Response and Security Teams (2020), FIRST). These best practices support the challenges of disclosing newly discovered vulnerabilities and, through different scenarios, describe the challenges and impacts for various stakeholders, i.e. finder, vendor, defenders and users. The VINCE platform (CERT/CC, [n. d.]) has been developed by CERT/CC and others to support multi-party disclosures and prevent many of the possible mistakes identified in the best-practice guide.
Moreover, many countries have introduced laws and regulations to stimulate stakeholders, with support from national CSIRTs, professional organisations, and others, all supporting the practice of vulnerability disclosure (Forum of Incident Response and Security Teams (2020), FIRST; oec, 2021; National Cyber Security Centre of The Netherlands (2018), NCSC-NL). Research institutes, universities, and companies such as Google and Facebook have adopted outgoing disclosure policies to protect the finders and make their intentions clear to receiving parties. (Reidsma et al., 2023; van der Horst, 2023).
A vulnerability disclosure to multiple parties is considered to be large-scale when there are five or more vendors involved and can become very complicated and stressful (Van Hove et al., 2023; Boucher and Anderson, 2022; Moura and Heidemann, 2023). However, vulnerability notification processes can quickly become large-scale, and the affected parties can be counted in the hundreds or even thousands. Identifying the contact information and notifying the responsible parties behind the vulnerable systems can be overwhelming for finders (Stock et al., 2018b; Çetin et al., 2017a; Maass et al., 2021; Stivala et al., 2024), which can eventually limit the development of vulnerability notification.
@@
Nonetheless, the current disclosure best practices are still struggling to catch up with the growing nature and importance of digital infrastructure. With best practices and prior experiences as a foundation for large-scale disclosure, the recent effort still suffers from the slow adoption of disclosure policy and incident handling among stakeholders. There are increasing large-scale internet-wide notification cases beyond the support of local or national security centres, which lead to finders having to carry out the notifications by setting up a messaging infrastructure themselves (Stivala et al., 2024; Chen et al., 2024). It has been reported that contact identification remains challenging, and the notification operation is still not broadly accepted among stakeholders (Moura and Heidemann, 2023; Chen et al., 2024). This shows that large-scale notification still requires new proposals to improve the process for both the notifying and receiving parties.
3. Problem Statement
How to disclose a vulnerability to a vendor is already well-considered, with CVD as the best practice, both for researchers (USENIX, 2024), as well as governments and vendors (oec, 2021). Over the last few years, more and more support mechanisms have come into place to support multi-party or even large-scale vulnerability disclosure. Still, it is unclear how well these are known in the academic security research community.
As we noted above, a large-scale vulnerability notification operation differs from disclosure and requires contacting other stakeholders differently. Publications often mention that it is challenging for security researchers, practitioners, and ethical hackers to reach out to the stakeholders in time. We examine the differences between disclosure and notification and examine the challenges associated with large-scale disclosure and notification operations.
From the above, we gather experiences from literature, community, and support organisations to propose best practices for the academic community to perform vulnerability disclosure and notification in the most efficient and effective way.
@@ 4. Literature Selection
We collected publications with extensive documentation on large-scale vulnerability disclosure or notification operations in the recent decade. We used literature search engines to find publications initially and performed a selection depending on the inclusion of disclosure procedures in stages and documentation before and after the disclosure or notification operations. We did not exhaustively locate all available disclosure or notification work. Rather, we selected work that can be representative of the large-scale disclosure and notification implementation in different disclosure and notification scenarios.
The initial search patterns we used were the combination of ”vulnerability”, ”disclosure”, ”notification”, ”large-scale”, and ”network” to look for the publications through major academic security conferences, literature databases, and search engines such as ACM Digital Library, IEEE Xplore, Scopus, and Google Scholar with all fields or metadata search filter enabled. We skimmed through the title, abstract, and partial content of search results that fit into the large-scale disclosure or notification scenarios discussed in Section 2.3. The initial result was broad and diverse. After a few iterations, we narrowed down our search keywords to ”vulnerability disclosure” and ”vulnerability notification”, with synonyms such as ”vulnerability alert” and ”vulnerability warning”. This gave us a short list of approximately 60 results that were relevant to vulnerability disclosure and notification.
To finalise the short list of publications, we examined the literature candidates with details such as assessment before an operation, selection of communication channels and messaging infrastructure, review after the operations, and contribution to best practices. We assigned different categories based on communication channels and involved stakeholders to seek the differences and choose work that has a measurable impact on the community. Furthermore, to understand the development of disclosure and notification best practices over the years, we aimed for publications that can be used to compare with each other, such as adopting, changing or improving methods based on other selected work. This means we looked into the reference list and related work sections of the publications to seek correlations and influences. Besides, we also selected work that made an impact in the cybersecurity community to illustrate the CVD as best practice in the last decade and to stimulate the discussion on how we can improve the existing guidelines. Finally, we separated the publications into vulnerability disclosure and notification operations as mentioned in Section 2.1.

### v2

@@ Abstract.
There has also been research into known vulnerabilities, using datasets or active scans to discover how many machines are still vulnerable. The ethical guidelines suggest that researchers also make an effort to notify the owners of these machines. We identify that this differs from vulnerability disclosure, but rather the practice of vulnerability notification. This practice has some similarities with vulnerability disclosure but should be distinguished from it, providing other challenges and requiring a different approach.
@@ 1. Introduction
Coordinated Vulnerability Disclosure (CVD), formerly known as responsible disclosure, is the best practice accepted in the security community for dealing with discovered vulnerabilities. When a new vulnerability is found, the finder conducts a risk and impact assessment and then initiates conversations with stakeholders to mitigate vulnerable systems or services in time. However, due to the shift in the network landscape, the scale of vulnerable systems and stakeholders involved has changed drastically from the past (van der Horst, 2023; Durumeric et al., 2014). The amount and variety of vulnerabilities have increased (Nakajima et al., 2019), and so have the affected parties (Boucher and Anderson, 2022; Li et al., 2016). This presents a new challenge for finders, particularly academic security researchers and practitioners: figuring out a scalable method to identify reliable contact information and ensure message delivery to stakeholders.
Over the years, support mechanisms, such as vulnerability disclosure policy (Reidsma et al., 2023), VINCE vulnerability platform (CERT/CC, [n. d.]), and bug bounty programs (Walshe and Simpson, 2020a), have come into place to handle the challenge in vulnerability disclosure. Finders and vendors can adopt CVD to perform disclosure with the support organisations such as Computer Security Incident Response Teams (CSIRTs) (CSI, [n. d.]) and Product Security Incident Response Teams (PSIRTs) (PSI, [n. d.]). However, not every finder or vendor has the same capacity and experience to conduct disclosure at scale (Walshe and Simpson, 2022b). Moreover, finders from different communities, such as academic security researchers (Boucher and Anderson, 2022), individual ethical hackers, practitioners, and bug bounty hunters, may have different security interests and constraints in selecting their approaches to reach stakeholders.
@@
Despite the challenge in vulnerability disclosure, notifications to stakeholders about known vulnerabilities and security issues that remain in existing systems have gained attention over the years (van der Horst, 2023). We identify the practice as vulnerability notification that informs end-users, such as product owners, hosting providers, domain owners, network operators, and incident responders. These stakeholders are different from vendors, who are traditionally the stakeholders in CVD. Whether or not the vulnerability is possibly known to the stakeholders and the public is the difference between vulnerability disclosure and vulnerability notification. In vulnerability disclosure, a finder discovers a new vulnerability and plans to disclose it to the vendor. Meanwhile, in vulnerability notification, a finder locates a known vulnerability on existing machines and notifies the end-users. This means that vulnerability notification is often conducted after the vulnerability disclosure. In particular, after a newly found vulnerability has been disclosed and mitigated with stakeholders or even made public, it is not guaranteed that all vulnerable systems are treated on a timely basis (Durumeric et al., 2014). Hence, vulnerability notification is there to improve the safety of the internet.
@@
Vulnerability notification is often initiated after a finder carries out network scanning (van der Ham and van Rijswijk-Deij, 2017) or vulnerability analysis on existing datasets (Chen et al., 2024) to locate the vulnerable systems with targeted vulnerabilities. A finder should retrieve the stakeholders’ contact information and select a communication channel to inform the affected parties. However, similar to the challenge in vulnerability disclosure, notification to stakeholders also suffers from the high number and complexity of stakeholders (Stivala et al., 2024; Çetin et al., 2017a). The contact retrieval and notification at scale pose an even tougher challenge to finders (Maass et al., 2021; Stock et al., 2016a) since stakeholders in vulnerability notification may have a more diverse and complex set of parties involved (Çetin et al., 2017a; Chen et al., 2024).
1.1. Research Questions
How to disclose a vulnerability to a vendor is well-considered, with CVD as the best practice and increasing support mechanisms suggested to finders, governments and vendors (oec, 2021). However, it is still unclear how well the best practices and mechanisms are known in the academic security research community. Furthermore, the lack of best practices for vulnerability notification has posed a growing challenge to not only academic security researchers but also other stakeholders. This motivates us to compile the research questions as follows:
  1. (1)
    What distinguishes vulnerability disclosure and notification, especially in communication to stakeholders?
  2. (2)
    How did academic researchers adopt best practices to carry out vulnerability disclosure and notification at scale?
  3. (3)
    What insights can we gain from the academic researchers’ experiences of large-scale disclosure and notification?
  4. (4)
    What are the best practices for researchers and other finders to perform vulnerability disclosure and notification?
In the following sections, we aim to answer each question with the insights we gain from literature, community, and support organisations, as well as our own disclosure and notification experience. In Section 2, we distinguish between vulnerability disclosure and notification, and explain how a large-scale scenario can affect the two distinct practices in operations. In Section 3, we explain how we select literature and perform a meta-review with our proposed stage model to extract experiences from the academic security research community. In Section 4, we compile the insights we gather from the academic experience in each stage of vulnerability disclosure and notification. Finally, in Section 5, we propose our best practices for the finders to address the current limitations of large-scale vulnerability disclosure and notification. The best practices can also help other stakeholders to improve their disclosure and notification handling.
@@ 2.1. Vulnerability Disclosure and Notification
Vulnerability disclosure, especially coordinated vulnerability disclosure, has been an accepted best practice in security research for years. There have been numerous publications and even standards published on this process (oec, 2021). However, notifying users with vulnerable systems has usually not been seen as a separate practice, but can come with challenges of its own. The difference between the two processes is illustrated in Figure 1.
In vulnerability disclosure, a finder discovers a new vulnerability and aims to share this with the vendor. The role of the finder in the following content includes both the responsibilities of identifying a vulnerability and informing the responsible parties (Forum of Incident Response and Security Teams (2020), FIRST). Once the disclosure process has started, the vulnerability is usually revealed for the first time. The finder and vendor(s) discuss the finding with or without a coordinator’s support. Subsequently, the vendor works on a mitigation. Vulnerability disclosure requires trusted channels and detailed message information to minimise disclosure leaks to unintended recipients. The finder may need to expect extra steps to communicate with the stakeholder on whether or not and how to disclose the vulnerabilities to the public, based on the disclosure policy and legal agreement on both ends. After some time, the vulnerability is made public, usually with a possible mitigation available.
In vulnerability notification, the vulnerability information is already available to the stakeholders before a notification starts. This means the vulnerability has already been documented and revealed to the stakeholders or the public. The weakness is likely documented with CVE numbers or discussed within specific communities. Potential malicious parties, such as criminals, may also be aware of the vulnerabilities and attempt to exploit them. The receiving users may or may not be aware of the vulnerability’s existence before seeing the notification message. Once they are aware of the issues, they may be able to find the vulnerability information with resources that are not limited to the finder. This can lead to different user response behaviours to the finder. The affected user may also have a different risk assessment based on the severity of the vulnerabilities. The aim of the notification process is to improve general security by removing vulnerable systems.
@@ 2.2. Evaluations of Disclosure and Notification
Vulnerability disclosure has received best practices over the years through the contributions from security researchers, vendors, and support organisations. There have been several publications describing the practices in detail and examining their effectiveness. Householder and Spring (Householder and Spring, 2022) proposed a model to assess the coordinator role in CVD, which provides insight into incident handling of vendors and the vulnerability lifecycle. Walshe and Simpson (Walshe and Simpson, 2022b) examined the effectiveness of how vendors with CVD programs and outsourced vulnerability platforms receive and process vulnerability disclosure. They revealed that disclosure program operators face a large number of vulnerability reports, which becomes a burden in vulnerability prioritisation. Nakajima et al. (Nakajima et al., 2019) looked into vulnerability management of IoT vendors across two countries and pointed out the disclosure pitfalls to avoid.
Vulnerability notification has received growing attention over the last decade, especially in the review of the notification efficiency. The Dutch Institute for Vulnerability Disclosure (DIVD) (div, [n. d.]) proposed the notification guideline (van der Horst, 2023) based on their framework and the Communication-Human Information Processing (C-HIP) model (Wogalter et al., 1999) with their notification operation experiences. Their contribution focuses on scalable notification strategies for end-users, such as an incident responder and an abuse specialist. Additionally, the challenges they documented with end-users resemble the capacity and awareness issues of program operators discussed by Walshe and Simpson in vulnerability disclosure(Walshe and Simpson, 2022b).
@@ 2.3. Large-Scale Disclosure and Notification
Large-scale disclosure operations have been reported as an increasing challenge by studies in the last decade (Householder and Spring, 2022; Walshe and Simpson, 2022b; Nakajima et al., 2019). Unlike the one-to-one or one-to-multiple disclosure to vendors, the complexity and number of stakeholders to inform have increased to a level beyond the effort individuals and small teams can make.
Many of the challenges associated with multi-party disclosure processes have been identified by CERT/CC (CERT/CC, 2024) and the FIRST Special Interest Group on Vulnerability Coordination (Vul, [n. d.]), resulting in a best practice document ”Guidelines and practices for Multi-Party Vulnerability Coordination and Disclosure” (Forum of Incident Response and Security Teams (2020), FIRST). These best practices address the challenges of disclosing newly discovered vulnerabilities and, through various scenarios, describe the challenges and impacts for different stakeholders, including finders, vendors, defenders, and users. The VINCE platform (CERT/CC, [n. d.]) has been developed by CERT/CC and other contributors to support multi-party disclosures and prevent many of the possible mistakes identified in the best practice guide.
Moreover, many countries have introduced laws, regulations (cra, [n. d.]), and vulnerability disclosure policies (BOD, 2020; National Cyber Security Centre of The Netherlands (2018), NCSC-NL) to stimulate stakeholders, with active support from national CSIRTs (CSI, [n. d.]) and PSIRTs (PSI, [n. d.]), professional organisations (oec, 2021; Forum of Incident Response and Security Teams (2020), FIRST), and others, all supporting the practice of vulnerability disclosure. Likewise, research institutes, universities, and companies, such as Google and Facebook, have adopted outgoing disclosure policies to protect the finders and make their intentions clear to receiving parties. (Reidsma et al., 2023; van der Horst, 2023).
A vulnerability disclosure to multiple parties is considered to be large-scale when there are five or more vendors involved, and can become very complicated and stressful (Van Hove et al., 2023; Boucher and Anderson, 2022; Moura and Heidemann, 2023). However, vulnerability notification processes can quickly become large-scale, and the affected parties can be counted in the hundreds or even thousands. Identifying the contact information and notifying the responsible parties behind the vulnerable systems can be overwhelming for finders (Stock et al., 2018b; Çetin et al., 2017a; Maass et al., 2021; Stivala et al., 2024), which can eventually limit the development of vulnerability notification.
@@
Nonetheless, the current best practices for vulnerability disclosure, especially vulnerability notification, are still struggling to catch up with the growing nature and importance of digital infrastructure. With best practices and prior experiences as a foundation for large-scale disclosure, the recent effort still suffers from the slow adoption of disclosure policy and incident handling among stakeholders (Van Hove et al., 2023; Chen et al., 2024). Although bug bounty programs, vulnerability disclosure policies, and regulations are adopted as best practices by the industry and countries, not every finder from different communities, such as academic security researcher, practitioner, ethical hacker, and bounty hunter, shares the same capacity and security interest to ensure the message delivery and conversations of the disclosure and notification at scale with the current best practices and guidelines (Walshe and Simpson, 2020a; Boucher and Anderson, 2022; Moura and Heidemann, 2023). For instance, academic security researchers may have time constraints for the publication schedule with their findings, while practitioners in a small team may have limited capacity and experience to prioritise stakeholders to inform when large-scale scenarios are considered. The current best practices do not necessarily cover the disclosure or notification guidelines for such finders.
There are increasingly large-scale internet-wide vulnerability notification cases beyond the support of local or national support organisations, which lead to finders, such as security researchers, carrying out the notifications by setting up a messaging infrastructure themselves (Stivala et al., 2024; Chen et al., 2024). It has been reported by finders that contact identification prioritisation remains challenging, and the notification operation is still not broadly accepted among stakeholders (Moura and Heidemann, 2023; Chen et al., 2024), even though disclosure policies or vulnerability report programs were present by stakeholders. This indicates that large-scale notifications still necessitate new proposals to enhance the process for both notifying and receiving parties.
So far, we have answered our first research question in Section 1.1. With a clear distinction between vulnerability disclosure and notification at scale, we aim to fill the gap by performing a meta-review on the experiences of academic security researchers, along with suggestions from other communities, support organisations, and our prior experience.
@@ 3. Literature Selection
To understand how the academic security research community approaches vulnerability disclosure and notification at scale over time, we collected publications with extensive documentation on large-scale vulnerability disclosure or notification operations in the last decade. We initially used literature search engines to find publications and then performed a selection based on the inclusion of disclosure procedures in stages and documentation before and after disclosure or notification operations. We did not exhaustively locate all available disclosure or notification work. Rather, we selected work that can be representative of the large-scale disclosure and notification implementation in different disclosure and notification scenarios from the academic security researcher’s perspective.
The initial search patterns we used were the combination of ”vulnerability”, ”disclosure”, ”notification”, ”report”, ”large-scale”, and ”network” to look for the publications through major academic security conferences, literature databases, and search engines such as ACM Digital Library, IEEE Xplore, Scopus, and Google Scholar with all fields or metadata search filter enabled. We skimmed through the title, abstract, and partial content of search results that fit into the large-scale disclosure or notification scenarios discussed in Section 2.3. The initial result was broad and diverse. After a few iterations, we narrowed down our search keywords to ”vulnerability disclosure” and ”vulnerability notification”, with synonyms such as ”vulnerability alert” and ”vulnerability warning”. It is worth noting that we also examine publications that don’t necessarily have such keywords but have related content with manual effort. This gave us a short list of approximately 60 results that were relevant to vulnerability disclosure and notification.
To finalise the short list of publications, we examined the literature candidates with details such as assessment before an operation, selection of communication channels and messaging infrastructure, review after the operations, and contribution to best practices. We assigned different categories based on communication channels and involved stakeholders to identify areas of difference and choose work that has a measurable impact on the stakeholders. Furthermore, to understand the development of disclosure and notification best practices over the years, we aimed for publications that can be used to compare with each other, such as adopting, changing or improving methods based on other selected work. This means we looked into the reference list and related work sections of the publications to seek correlations and influences. Besides, we also selected work that made an impact in the security communities to illustrate the CVD as best practice and to stimulate the discussion on how we can improve the existing guidelines. Finally, we separated the publications into vulnerability disclosure and notification operations as mentioned in Section 2.1.